# Well-posedness of McKean-Vlasov generalized multivalued BSDEs

### v1

@@ 3. BSVIs with locally Lipschitz and non-linear growth conditions
$\quad+\frac{C}{N^{2(1-\alpha)}}[\mathcal{M}(T)+\mathcal{M}^{2}(T)].$
By Grownwall inequality, we have
$\mathbb{E}\sup_{0\leq s\leq T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\leq e^{(5+10r_{N}^{2})T}\big\{\rho_{N}(F_{n}-F)+\rho_{N}(F_{m}-F)+\frac{C}{N^{2(1-\alpha)}}[\mathcal{M}(T)+\mathcal{M}^{2}(T)]\big\}.$ (19)
It follows that
$\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\mathrm{d}s$ (20)
@@
If $r_{N}^{2}\leq\frac{1-\alpha}{5T}logN-\frac{1}{2}$, then by passing to the limit in $n,m,N$, we deduce that $(Y^{\frac{1}{n}},Z^{\frac{1}{n}})$ forms a Cauchy sequence in the Banach space $(\mathcal{S}_{k}^{\lambda,\mu}\cap\mathcal{H}_{k}^{\lambda,\mu}\cap\widetilde{\mathcal{H}}_{k}^{\lambda,\mu})\times\mathcal{H}_{k\times d}^{\lambda,\mu}$. Hence, there exists $(Y,Z)$ such that
$\lim_{n\to\infty}\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}|^{2}\mathrm{d}s=0,\quad\lim_{n\to\infty}\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}|^{2}\mathrm{d}A_{s}=0,\qquad$ (22)
$\lim_{n\to\infty}\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\|Z_{s}^{\frac{1}{n}}-Z_{s}\|^{2}\mathrm{d}s=0,\quad\lim_{n\to\infty}\mathbb{E}\sup_{0\leq s\leq T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}|^{2}=0.\qquad$
For any given $N$, we define $H_{n}^{N}:=\{(s,w):|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}|^{2}\geq N^{2}\}$ and $\bar{H}_{n}^{N}=\Omega-H_{n}^{N}$. By an argument analogous to the estimations of $\Gamma_{1}$ and $\Gamma_{3}$, we have
[…]
Passing to the limit in $n$ and $N$, we obtain
@@
Step 4): The proof of uniqueness follows the same method as in Proposition 2 and is therefore omitted.

### v2

@@ 3. BSVIs with locally Lipschitz and non-linear growth conditions
$\quad+C\mathcal{M}(T)\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\dfrac{|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2}}{N^{2}}1_{D_{n,m}^{N}}\mathrm{d}s$
$\quad+\frac{40L^{2}}{N^{2}}\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+\|Z_{s}^{\frac{1}{n}}\|^{2}+\|Z_{s}^{\frac{1}{m}}\|^{2}1_{D_{n,m}^{N}}\mathrm{d}s$
$\quad+\frac{10L^{2}}{N^{2}}\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big(|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+\|Z_{s}^{\frac{1}{n}}\|^{2}+\|Z_{s}^{\frac{1}{m}}\|^{2}\big)^{2}1_{D_{n,m}^{N}}\mathrm{d}s.$
Applying Itô’s formula to $e^{\lambda s+\mu A_{s}}|Y_{s}|^{4}$, and employing Hölder’s inequality along with similar techniques as in the proof of (18), we can obtain results with a similar form but different orders. Since the derivation is rather lengthy, we omit the details here. Denoting the corresponding upper bound on the right-hand side as $\tilde{\mathcal{M}}(T)$, we have
$\mathbb{E}[\Gamma_{1}]\leq\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}s+\frac{C}{N^{2}}\tilde{\mathcal{M}(T)}+\frac{C}{N^{2}}\mathcal{M}^{2}(T)+\frac{C}{N^{2}}\mathcal{M}(T).$
Other estimates are as follows:
$\Gamma_{2}+\Gamma_{4}\leq 2\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}s$
$\quad+\int_{t}^{T}e^{\lambda s+\mu A_{s}}|F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))|^{2}1_{\bar{D}_{n,m}^{N}}\mathrm{d}s$
$\quad+\int_{t}^{T}e^{\lambda s+\mu A_{s}}|F(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))|^{2}1_{\bar{D}_{n,m}^{N}}\mathrm{d}s,$
$\Gamma_{5}\leq(K^{2}+1)\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}A_{s},$
@@
$\quad+\frac{1}{4}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big[\mathbb{E}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}+\mathbb{E}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\big]1_{\bar{D}_{n,m}^{N}}\mathrm{d}s.$
Using the similar proof to Proposition 2, we have
$e^{\lambda t+\mu A_{t}}|Y_{t}^{\frac{1}{n}}-Y_{t}^{\frac{1}{m}}|^{2}+\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}(\lambda\mathrm{d}s+\mu\mathrm{d}A_{s})+\int_{t}^{T}e^{\lambda s+\mu A_{s}}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\mathrm{d}s$
$\quad+2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\nabla\varphi_{n}(Y_{s}^{\frac{1}{n}})-\nabla\varphi_{m}(Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}s$
$\quad+2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},\nabla\psi_{n}(Y_{s}^{\frac{1}{n}})-\nabla\psi_{m}(Y_{s}^{\frac{1}{m}})\rangle\mathrm{d}A_{s}$
[…]
$\quad+\frac{1}{4}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big[\mathbb{E}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}+\mathbb{E}\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\big]1_{\bar{D}_{n,m}^{N}}\mathrm{d}s$
[…]
$\quad+C\mathcal{M}(T)\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\dfrac{|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}^{\frac{1}{m}}|^{2}}{N^{2}}1_{D_{n,m}^{N}}\mathrm{d}s$
$\quad+\frac{40L^{2}}{N^{2}}\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+\|Z_{s}^{\frac{1}{n}}\|^{2}+\|Z_{s}^{\frac{1}{m}}\|^{2}1_{D_{n,m}^{N}}\mathrm{d}s$
$\quad+\frac{10L^{2}}{N^{2}}\mathbb{E}\int_{t}^{T}e^{\lambda s+\mu A_{s}}\big(|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}^{\frac{1}{m}}|^{2}+\|Z_{s}^{\frac{1}{n}}\|^{2}+\|Z_{s}^{\frac{1}{m}}\|^{2}\big)^{2}1_{D_{n,m}^{N}}\mathrm{d}s$
$\quad+\int_{t}^{T}e^{\lambda s+\mu A_{s}}|F_{n}(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))-F(s,Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}},\mathcal{L}(Y_{s}^{\frac{1}{n}},Z_{s}^{\frac{1}{n}}))|^{2}1_{\bar{D}_{n,m}^{N}}\mathrm{d}s$
$\quad+\int_{t}^{T}e^{\lambda s+\mu A_{s}}|F(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))-F_{m}(s,Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}},\mathcal{L}(Y_{s}^{\frac{1}{m}},Z_{s}^{\frac{1}{m}}))|^{2}1_{\bar{D}_{n,m}^{N}}\mathrm{d}s$
$\quad+(K^{2}+1)\int_{t}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}\mathrm{d}A_{s}$
$\quad-2\int_{t}^{T}e^{\lambda s+\mu A_{s}}\langle\langle Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}},(Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}})dW_{s}\rangle.$
By taking the expectation on both sides of the previous estimate at $t=0$, and considering the constraints on $\lambda,\mu$ as well as Lemma 8, it follows that
$\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\big(|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}(\mathrm{d}s+\mathrm{d}A_{s})+\|Z_{s}^{\frac{1}{n}}-Z_{s}^{\frac{1}{m}}\|^{2}\mathrm{d}s\big)$ (19)
$\leq\frac{8}{5}e^{\frac{192Tr_{N}^{2}}{5}}\big\{\rho_{N}(F_{n}-F)+\rho_{N}(F_{m}-F)+\frac{C}{N^{2}}\tilde{\mathcal{M}(T)}+\frac{C}{N^{2}}\mathcal{M}^{2}(T)+\frac{C}{N^{2}}\mathcal{M}(T)\big\}.\quad$
Taking the supremum over time and applying the BDG inequality, we obtain
$\mathbb{E}\sup_{0\leq s\leq T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}^{\frac{1}{m}}|^{2}$ (20)
$\leq[\frac{4}{3}+(C+\frac{256}{5}r_{N}^{2})e^{\frac{192Tr_{N}^{2}}{5}}]\big\{\rho_{N}(F_{n}-F)+\rho_{N}(F_{m}-F)+\frac{C}{N^{2}}\tilde{\mathcal{M}(T)}+\frac{C}{N^{2}}\mathcal{M}^{2}(T)+\frac{C}{N^{2}}\mathcal{M}(T)\big\}.$
@@
$\leq[\frac{4}{3}+(C+\frac{256}{5}r_{N}^{2})e^{\frac{192Tr_{N}^{2}}{5}}]\big\{\rho_{N}(F_{n}-F)+\rho_{N}(F_{m}-F)+\frac{C}{N^{2}}\tilde{\mathcal{M}(T)}+\frac{C}{N^{2}}\mathcal{M}^{2}(T)+\frac{C}{N^{2}}\mathcal{M}(T)\big\}$
$\leq[\frac{4}{3}+(C+\frac{256}{5}r_{N}^{2})e^{\frac{192Tr_{N}^{2}}{5}}]\big\{\rho_{N}(F_{n}-F)+\rho_{N}(F_{m}-F)+\frac{C}{N^{2(1-\alpha)}}[\tilde{\mathcal{M}(T)}+\mathcal{M}^{2}(T)+\mathcal{M}(T)]\big\}.$
If $r_{N}^{2}\leq\dfrac{5(1-\alpha)}{96T}logN$, then by passing to the limit in $n,m,N$, we deduce that $(Y^{\frac{1}{n}},Z^{\frac{1}{n}})$ forms a Cauchy sequence in the Banach space $(\mathcal{S}_{k}^{\lambda,\mu}\cap\mathcal{H}_{k}^{\lambda,\mu}\cap\widetilde{\mathcal{H}}_{k}^{\lambda,\mu})\times\mathcal{H}_{k\times d}^{\lambda,\mu}$. Hence, there exists $(Y,Z)$ such that
$\lim_{n\to\infty}\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}|^{2}\mathrm{d}s=0,\quad\lim_{n\to\infty}\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}|^{2}\mathrm{d}A_{s}=0,$ (22)
$\lim_{n\to\infty}\mathbb{E}\int_{0}^{T}e^{\lambda s+\mu A_{s}}\|Z_{s}^{\frac{1}{n}}-Z_{s}\|^{2}\mathrm{d}s=0,\quad\lim_{n\to\infty}\mathbb{E}\sup_{0\leq s\leq T}e^{\lambda s+\mu A_{s}}|Y_{s}^{\frac{1}{n}}-Y_{s}|^{2}=0.$
For any given NN, we define $H_{n}^{N}:=\{(s,w):|Y_{s}^{\frac{1}{n}}|^{2}+|Y_{s}|^{2}+|Z_{s}^{\frac{1}{n}}|^{2}+|Z_{s}|^{2}\geq N^{2}\}$ and $\bar{H}_{n}^{N}=\Omega-H_{n}^{N}$.By an argument analogous to the estimations of $\Gamma_{1}$ and $\Gamma_{3}$, we have
$\leq\frac{C}{N^{2}}\tilde{\mathcal{M}(T)}+\frac{C}{N^{2}}\mathcal{M}^{2}(T)+\frac{C}{N^{2}}\mathcal{M}(T)+2\rho_{N}(F_{n}-F)$
@@
Passing to the limit in $n$ and NN , we obtain
@@
The proof of uniqueness follows the same method as in Proposition 2 and is therefore omitted.